difference between a man age and his son age is 23 years and their sum is 41 years. find the age of the man

Answer by checkley79(3341)

X-Y=23
X+Y=41 ADD.
-----------------
2X=64
X=64/2
X=32 ANS. FOR THE MAN.
32-Y=23
-Y=23-32
-Y=-9
Y=9 ANS. FOR THE SON.
PROOF:
32+9=41
41=41